OpenStudy (anonymous):

Rex was building a rectangular porch. Laying boards end to end for one length of the porch, he found that eight boards were four feet too long and seven boards were two feet too short. If the porch is 9 ft. wide, what is the area of the porch?

OpenStudy (anonymous):

given : w=9 l= (let's make an equation base on the given) =8(x-4) + 7(x+2) solve first for the length. l=8(x-4) + 7(x+2) 8(x-4) + 7(x+2)=0 8x-32+ 7x+14 =0 x-18 =0 x =18 substitute the value to find the length. l= 8(18-4) + 7(18+2) l=252 to find the area. A=lw =252(9) therefor: A(area)=2268 ft.